# Diverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp-core enigma

### v1

@@ 3.5 Driving forces of core formation
Fig. 5 shows the central dark matter densities at $r=500$ pc as a function of cosmic time. We choose this radius as it is small enough to showcase the differences between the models explored whilst being large enough to ensure we are not being dominated by numerical heating effects, i.e. we choose a radius that is at least the size of the Power radius for all set-ups explored. Again we average the dark matter densities over time bins of $\Delta T=500$ Myr and indicate the mean densities as solid lines and the standard deviation in each bin by the shaded regions.
@@ 3.5.1 SN feedback
This highly variable behaviour is quite clearly illustrated by the m10q set-up in the first column. With CRs, the system experiences a higher initial burst which is powerful enough to quench the dwarf system from $z\sim 6$. Hence both m10qCR and m10qCR+AGN end up with cuspy profiles, even though the cumulative injected SN energy for m10qCR+AGN slighly exceeds the $z=0$ binding energy, as the absence of late-time star formation means that a cusp can be re-established quite easily after the initial burst. Note that m10qCR has a slighly reduced central density at $z=0$ compared to m10qCR+AGN due to low-level late-time star formation activity; however, these bursts are not powerful enough to induce a core. Without CRs, the initial star formation burst is less powerful, allowing for continued star formation activity. For the m10q set-up, there are a few more powerful bursts at intermediate redshifts, which then quench the systems at much later times from $z\sim 1.5$. For the m10qAGN set-up, the AGN regulation means that the intermediate redshift star formation bursts are less intense so that star formation can continue until $z=0$. Both m10q and m10qAGN end up with cumulative SN energy almost an order of magnitude above the $z=0$ binding energy leading to cored profiles; however, m10qAGN has a more extreme core due to the continued star formation activity (also see Muni et al., 2024).
@@ 3.5.2 AGN feedback
In the fire-3 model, AGN feedback is injected in several channels including mechanical feedback, radiative feedback and CRs (see Section 2.2.1 for details). Comparing this multi-channel AGN energy to the SN energy that is predominantly injected mechanically is not straightforward, especially since we are most concerned with feedback-induced gravitational potential perturbations. Therefore, we need to assess how the AGN energy will couple dynamically. Fully tracking and analysing coupling efficiencies of different AGN feedback modes is beyond the scope of this paper, so we make some simplifying assumptions based on previous literature.
@@ 3.6 Cosmic evolution of central dark matter densities in observational context
The first panel demonstrates that our simulations are generally in good agreement with the observed central densities of nearby dwarfs, keeping in mind that the sample sizes are limited in both cases (16 observed dwarfs versus 12 simulated dwarfs). Two of the m10z simulations, m10zCR and m10zAGN, lead to very strong cores, with the central densities being more severely suppressed than in the observations. Our simulations do not reproduce the observed strong cusps with central densities above the NFW mean relation. However, we note that the vast majority of our dwarfs have late star formation histories, with only the CR set-ups for m10q and m10y resulting in an early truncation of star formation, as defined by Read et al. (2019), with no activity for the past 6 Gyr. These dwarfs with early truncation are associated with cuspy profiles hinting that we would require more early-forming set-ups to reproduce these strong cores.
@@ 4.1 Comparison with previous theoretical work
Previous studies found systematic yet small-scale suppressions of central dark matter densities by AGN feedback in dwarfs (Koudmani et al., 2022; Arjona-Galvez et al., 2024), yet in both cases, the simulations did not explicitly resolve the multiphase ISM. Cusp-to-core transformations with AGN feedback were also explored by Arora et al. (2024). However, they employed the fiducial Bondi accretion model for BH growth, which suppresses the growth of low-mass BHs such that AGN activity in dwarfs is insignificant in these simulations and does not affect dark matter distributions.
In our study, we investigate, for the first time, the impact of AGN feedback on dark matter profiles in dwarfs with a resolved multi-phase ISM model and a BH growth scheme that does not suppress AGN activity in dwarf galaxies. Several works exploring BH physics within the fire model have found that CRs appear to be a key ingredient for effective AGN feedback (e.g. Su et al., 2021; Wellons et al., 2023; Byrne et al., 2023b), so we also investigate the impact of CR feedback on the interplay between baryons and dark matter in dwarfs. We find that efficient AGN feedback can actively drive core formation and, together with CRs, significantly enhances the diversity of dwarf rotation curves by leading to more varied star formation histories. Indeed, in many cases our runs including CRs (with or without AGN) lead to cuspier profiles due to suppressed star formation activity. This is in contrast with recent findings from Martin-Alvarez et al. (2023), who find that CRs are prone to enhance core formation. We note that for one of our set-ups (compare m10z to m10zCR), we also observe this behaviour, and indeed it is plausible that CRs could act to either suppress or enhance cores depending on their interaction with the ISM and the assembly history of the given dwarf galaxy. Our most important finding is that BHs and CRs as additional baryonic processes lead to varying overall feedback efficiencies that can lead to both cuspier or more cored profiles for a given assembly history.
@@ 4.2 Observational uncertainties
Apart from the observational uncertainties in inferring dark matter mass distributions, it also remains extremely challenging to determine the drivers behind core formation from observations. In particular, there are no observational constraints on the potential role of AGN in cusp-to-core transformations. This is mainly due to the number of observed AGN in nearby dwarf galaxies being very limited. To further complicate matters, AGN are associated with disturbed gas rotation curves – see e.g. Manzano-King & Canalizo (2020), who find a strong association between AGN activity and disturbed gas kinematics in observed dwarf rotation curves, making it even more difficult to establish a link between AGN activity and central dark matter densities. JWST may partly alleviate these issues by disentangling rotational and outflow components with high-resolution NIRSpec-IFU observations (e.g. Bohn et al., 2023). From our simulations, we predict that if cored dwarf galaxies were found below the critical mass ratio of $\left(\frac{M_{\mathrm{stellar}}}{M_{\mathrm{vir}}}\right)_{\mathrm{crit}}=5\times 10^{4}$ this would be a strong indicator of AGN-driven core formation in dwarfs.
@@ Acknowledgements
The authors are grateful for helpful discussions with Vasily Belokurov, Martin Bourne, Jenny Greene, Vid Irsic and Sergio Martin-Alvarez. The authors would also like to thank Isabel Santos-Santos for providing the observational rotation curve data. The simulations presented in this work were run on the Flatiron Institute’s research computing facilities (the Iron compute cluster), supported by the Simons Foundation. SK has been supported by a Flatiron Research Fellowship and a Junior Research Fellowship from St Catharine’s College, Cambridge. The Flatiron Institute is supported by the Simons Foundation. DAA acknowledges support by NSF grant AST-2108944, NASA grant ATP23-0156, STScI grants JWST-GO-01712.009-A and JWST-AR-04357.001-A, Simons Foundation Award CCA-1018464, and Cottrell Scholar Award CS-CSA-2023-028 by the Research Corporation for Science Advancement. Support for ISS was provided by NSF Collaborative Research Grant 2108318. SW received support from the NASA RIA grant 80NSSC24K0838.

### v2

@@ 3.5 Driving forces of core formation
Fig. 6 shows the central dark matter densities at $r=150$ pc (top panel) and $r=500$ pc (bottom panel) as a function of cosmic time. In addition to our fiducial $r=150$ pc radius, an important reference point for observations, we also present the density evolution at $r=500$ pc which is at least the size of the Power radius for all set-ups explored. Again we average the dark matter densities over time bins of $\Delta T=500$ Myr and indicate the mean densities as solid and dotted lines and the standard deviation in each bin by the shaded regions.
@@ 3.5.1 SN feedback
This highly variable behaviour is quite clearly illustrated by the m10q set-up in the first column. With CRs, the system experiences a higher initial burst which is powerful enough to quench the dwarf system from $z\sim 6$. Hence both m10qCR and m10qCR+AGN end up with cuspy profiles, even though the cumulative injected SN energy for m10qCR+AGN slighly exceeds the $z=0$ binding energy, as the absence of late-time star formation means that a cusp can be re-established quite easily after the initial burst. Note that m10qCR has a slighly reduced central density at $z=0$ compared to m10qCR+AGN due to low-level late-time star formation activity; however, these bursts are not powerful enough to induce a core. Without CRs, the initial star formation burst is less powerful, allowing for continued star formation activity. For the m10q set-up, there are a few more powerful bursts at intermediate redshifts, which then quench the systems at much later times from $z\sim 1.5$. For the m10qAGN set-up, the AGN regulation means that the intermediate redshift star formation bursts are less intense so that star formation can continue until $z=0$. Both m10q and m10qAGN end up with cumulative SN energy almost an order of magnitude above the $z=0$ binding energy leading to cored profiles; however, m10qAGN has a more extreme core due to the continued star formation activity (also see Muni et al., 2025).
@@ 3.5.2 AGN feedback
In the fire-3 model, AGN feedback is injected in several channels including mechanical feedback, radiative feedback and CRs (see Section 2.2.1 for details). Comparing this multi-channel AGN energy to the SN energy that is predominantly injected mechanically is not straightforward, especially since we are most concerned with feedback-induced gravitational potential perturbations. Therefore, we need to assess how the AGN energy will couple dynamically. Fully tracking and analysing coupling efficiencies of different AGN feedback modes is beyond the scope of this paper, so we make some simplifying assumptions based on the previous literature.
@@ 3.6 Cosmic evolution of central dark matter densities in observational context
The first panel demonstrates that our simulations are generally in good agreement with the observed central densities of nearby dwarfs, keeping in mind that the sample sizes are limited in both cases (16 observed dwarfs versus 12 simulated dwarfs). Two of the m10z simulations, m10zCR and m10zAGN, lead to very strong cores, with the central densities being more severely suppressed than in the observations. Our simulations do not reproduce the observed strong cusps with central densities above the NFW mean relation. However, we note that the majority of our dwarfs have late star formation histories, with only the fiducial and CR set-ups for m10q and CR set-ups for m10y resulting in an early truncation of star formation, as defined by Read et al. (2019), with no (significant) activity for the past 6 Gyr. These dwarfs with early truncation are associated with cuspier profiles hinting that we would require more early-forming set-ups to reproduce these strong cusps.
@@ 4.1 Comparison with previous theoretical work
Previous studies found systematic yet small-scale suppressions of central dark matter densities by AGN feedback in dwarfs (Koudmani et al., 2022; Arjona-Gálvez et al., 2024), yet in both cases, the simulations did not explicitly resolve the multiphase ISM. Cusp-to-core transformations with AGN feedback were also explored by Arora et al. (2024). However, they employed the fiducial Bondi accretion model for BH growth, which suppresses the growth of low-mass BHs such that AGN activity in dwarfs is insignificant in these simulations and does not affect dark matter distributions.
In our study, we investigate, for the first time, the impact of AGN feedback on dark matter profiles in dwarfs with a resolved multi-phase ISM model and a BH growth scheme that does not suppress AGN activity in dwarf galaxies. Several works exploring BH physics within the fire model have found that CRs appear to be a key ingredient for effective AGN feedback (e.g. Su et al., 2021; Wellons et al., 2023; Byrne et al., 2024), so we also investigate the impact of CR feedback on the interplay between baryons and dark matter in dwarfs. We find that efficient AGN feedback can actively drive core formation and, together with CRs, significantly enhances the diversity of dwarf rotation curves by leading to more varied star formation histories. Indeed, in many cases our runs including CRs (with or without AGN) lead to cuspier profiles due to suppressed star formation activity. This is in contrast with recent findings from Martin-Alvarez et al. (2023), who find that CRs are prone to enhance core formation. We note that for one of our set-ups (compare m10z to m10zCR), we also observe this behaviour, and indeed it is plausible that CRs could act to either suppress or enhance cores depending on their interaction with the ISM and the assembly history of the given dwarf galaxy. Our most important finding is that BHs and CRs as additional baryonic processes lead to varying overall feedback efficiencies that can lead to both cuspier or more cored profiles for a given assembly history.
@@ 4.2 Observational uncertainties
Apart from the observational uncertainties in inferring dark matter mass distributions, it also remains extremely challenging to determine the drivers behind core formation from observations. In particular, there are no observational constraints on the potential role of AGN in cusp-to-core transformations. This is mainly due to the number of observed AGN in nearby dwarf galaxies being very limited. To further complicate matters, AGN are associated with disturbed gas rotation curves – see e.g. Manzano-King & Canalizo (2020), who find a strong association between AGN activity and disturbed gas kinematics in observed dwarf rotation curves, making it even more difficult to establish a link between AGN activity and central dark matter densities. Current and upcoming IFU surveys may partly alleviate these issues by disentangling rotational and outflow components, e.g. using MaNGA (see e.g. Rodríguez Morales et al., 2025) or high-resolution NIRSpec-IFU observations with JWST (e.g. Bohn et al., 2023). From our simulations, we predict that if cored dwarf galaxies were found below the critical mass ratio of $\left(\frac{M_{\mathrm{stellar}}}{M_{\mathrm{vir}}}\right)_{\mathrm{crit}}=5\times 10^{4}$ this would be a strong indicator of AGN-driven core formation in dwarfs.
@@ Acknowledgements
The authors would like to thank the anonymous referee whose comments improved the quality of this paper. The authors are grateful for helpful discussions with Vasily Belokurov, Martin Bourne, Jenny Greene, Vid Irsic and Sergio Martin-Alvarez. The authors would also like to thank Isabel Santos-Santos for providing the observational rotation curve data. The simulations presented in this work were run on the Flatiron Institute’s research computing facilities (the Iron compute cluster), supported by the Simons Foundation. SK has been supported by a Flatiron Research Fellowship, a Junior Research Fellowship from St Catharine’s College, Cambridge and a Research Fellowship from the Royal Commission for the Exhibition of 1851. The Flatiron Institute is supported by the Simons Foundation. DAA acknowledges support by NSF grant AST-2108944, NASA grant ATP23-0156, STScI grants JWST-GO-01712.009-A and JWST-AR-04357.001-A, Simons Foundation Award CCA-1018464, and Cottrell Scholar Award CS-CSA-2023-028 by the Research Corporation for Science Advancement. Support for ISS was provided by NSF Collaborative Research Grant 2108318. SW received support from the NASA RIA grant 80NSSC24K0838.